# Magnon Correlation Enables Spin Injection, Dephasing, and Transport in Canted Antiferromagnets

### v1

@@ Abstract
Figure 2: (a) Two-sublattice model of AFM. The dashed square represents a unit cell with a lattice constant $a$. (b) and (c) Polarizations ${\bf m}$ by green arrows of the anti-phase mode “$1$” and in-phase mode “$2$” with the Néel vector ${\bf n}$ of mode “1” (“2”) oscillating along the $\hat{\bf y}~{}(\hat{\bf x})$-axis, indicated by orange arrows. (d) Band structure of two modes. Inset: average frequency difference $|\delta\omega|=|\omega_{1}-\omega_{2}|$ over $|{\bf k}|\in[0,0.5]$ nm-1.
@@
According to the spin matrix (1), the total spin operator along the Néel vector
$\hat{S}_{z}=\sum_{{\bf k}}\left({\cal A}_{\bf k}\hat{b}_{1,{\bf k}}^{\dagger}\hat{b}_{2,{\bf k}}+{\cal B}_{\bf k}\hat{b}_{1,{\bf k}}^{\dagger}\hat{b}^{\dagger}_{2,{-\bf k}}\right)+{\rm H.c.},$ (3)
@@
with expressions for ${\cal A}_{\bf k}$ and ${\cal B}_{\bf k}$ with arbitrary canted angles $\theta$ given in SM [50]. The population of both magnon modes carries no spin along the Néel vector; the spin is solely carried by the correlations between the two modes. A solution ${\cal A}_{\bf k}\approx 1$ and ${\cal B}_{\bf k}\approx 0$ when the canted angle is small with $|\gamma|\hbar\mu_{0}H\ll 8JS$ is sufficient to analyze the mechanism. The spin current and torque for spin polarization along ${\bf n}$ are also solely carried by magnon correlations. Indeed, according to Eqs. (2a) and (2b),
@@
without contribution from the magnon population. Due to the breaking of $U(1)$ symmetry in canted AFMs, $S_{z}$ is no longer conserved by a spin torque $\hat{T}_{0z}$.
@@
Electrical injection of magnon correlation.—We examine the electrical spin injection from heavy metals to canted AFMs in a bilayer structure as illustrated in Fig. 1(b). When a charge current is applied along the $\hat{\bf y}$-direction in the metals, the injected spin current along $\hat{\bf x}$ generates a spin accumulation $\mu_{\uparrow}\neq\mu_{\downarrow}$ polarized along the Néel vector $\hat{\bf z}$ at the interface through the spin Hall effect. This accumulated spin of electrons then excites magnon correlations in the underlying AFM as in Fig. 1(b) via the $s$-$d$ exchange interaction between local spins and conduction electrons at the interface, governed by the Hamiltonian $\hat{H}_{\rm int}=\sum_{i\in{\rm int}}^{N_{\rm int}}\sum_{l=1}^{\cal N}{\cal J}_{l}\hat{\bf s}_{i}\cdot\hat{\bf S}_{li}$ [56, 57, 58], where the summation is performed over the magnetic unit cells “$i$”, with $N_{\rm int}$ being their total number in the interfacial plane, and the spin sublattices “$l$”. ${\cal J}_{l}$ represents the local exchange coupling, which typically depends on the AFM sublattices. Equal or unequal coupling with different sublattices corresponds to compensated and uncompensated interfaces, as observed experimentally [59, 38, 45].
@@
For magnons in the two-sublattice model with a lattice constant $a$ [Fig. 2(a)] interacting with the electrons in a heavy metal of thickness $d$ [50, 56, 57, 58],
@@
×$\times\hat{c}_{{{\bf k}^{\prime}},\alpha}^{\dagger}\hat{c}_{{{\bf k}},\beta}\hat{b}_{\zeta,{\bf q}}\delta_{{{\bf k}}+{\bf q},{\bf k}^{\prime}}+{\rm H.c.},$ (5)
@@
are governed by exchange couplings ${\cal J}_{1,2}$ and polarizations of sublattice spins $\delta{\bf S}_{1,{\bf q}}^{(1)}=(\sqrt{2}/2)(-B_{1,\bf q}\cos\theta\hat{\bf x}+iA_{1,\bf q}\hat{\bf y}+B_{1,\bf q}\sin\theta\hat{\bf z})$ and $\delta{\bf S}_{2,{\bf q}}^{(1)}=(\sqrt{2}/2)(-B_{1,\bf q}\cos\theta\hat{\bf x}-iA_{1,\bf q}\hat{\bf y}-B_{1,\bf q}\sin\theta\hat{\bf z})$ in mode “$1$” and $\delta{\bf S}_{1,{\bf q}}^{(2)}=(\sqrt{2}/2)(A_{2,{\bf q}}\cos\theta\hat{\bf x}-iB_{2,{\bf q}}\hat{\bf y}-A_{2,{\bf q}}\sin\theta\hat{\bf z})$ and $\delta{\bf S}_{2,{\bf q}}^{(2)}=(\sqrt{2}/2)(-A_{2,{\bf q}}\cos\theta\hat{\bf x}-iB_{2,{\bf q}}\hat{\bf y}-A_{2,{\bf q}}\sin\theta\hat{\bf z})$ in mode “$2$”, as shown in Fig. 2(b) and (c). When the exchange interaction $8JS\gg\hbar\omega_{1,2}({\bf q})$ dominates over the other energies, $A_{1,2}({\bf q})\approx\sqrt{8JS/\hbar\omega_{1,2}({\bf q})}$ and $B_{1,2}({\bf q})\approx 1/A_{1,2}({\bf q})$, and, thereby, $A_{1,2}({\bf q})\gg B_{1,2}({\bf q})$. When ${\cal J}_{1}={\cal J}_{2}$, $|\boldsymbol{\chi}^{(1),(2)}_{\bf q}|\propto B_{1,2}({\bf q})$; when ${\cal J}_{1}\gg{\cal J}_{2}$ or ${\cal J}_{2}\gg{\cal J}_{1}$, $|\boldsymbol{\chi}^{(1),(2)}_{\bf q}|\propto A_{1,2}({\bf q})$, so the electron-magnon coupling is enhanced when ${\cal J}_{1}\neq{\cal J}_{2}$.
$|\boldsymbol{\sigma}_{\alpha\beta}\cdot\boldsymbol{\chi}^{(\zeta)}_{\bf q}|^{2}$ is responsible for the excitation of magnon population in the $\zeta$-mode; $(\boldsymbol{\sigma}_{\alpha\beta}\cdot\boldsymbol{\chi}^{(1)}_{\bf q})(\boldsymbol{\sigma}_{\beta\alpha}\cdot\boldsymbol{\chi}^{(2)\ast}_{\bf q})$ denotes the mixing excitation of both modes in a net spin-flip process, which creates a magnon correction as in Fig. 1(b). The latter process vanishes in collinear AFMs. Since the magnon population carries no spin along $\hat{\bf z}$, the electron spin is not conserved in the spin-flip scattering (5) since $\boldsymbol{\chi}^{(1,2)}_{\bf q}\nparallel\hat{\bf z}$. When the canted angle $\theta$ is small, $\chi_{z}\rightarrow 0$ such that the interfacial spin-conserving scattering $|\boldsymbol{\sigma}_{\alpha\alpha}\cdot\boldsymbol{\chi}^{(\zeta)}_{\bf q}|^{2}\rightarrow 0$. Consequently, the excitation of magnon correlation is dominated by electron spin-flip scattering.
@@
$\partial_{t}{\rho}_{\bf q}|_{\rm inj}$ $=-\frac{\pi}{\hbar}\frac{S}{2N_{\rm int}}\frac{a^{2}}{d^{2}}\sum_{{\bf k},{\bf k}^{\prime}}\sum_{\alpha\beta}\left[{\rho}_{\bf q}-\sigma_{3}N_{0}(\varepsilon_{{\bf k}^{\prime},\alpha}-\varepsilon_{{\bf k},\beta})\right]$
×$\times(f_{{\bf k},\beta}-f_{{\bf k}^{\prime},\alpha}){\cal M}_{\alpha\beta}({\bf k},{\bf k}^{\prime})\delta_{{\bf k}+{\bf q},{\bf k}^{\prime}}+{\rm H.c.},$ (7)
where $N_{0}(\varepsilon)=\{\exp[{\varepsilon}/(k_{B}T)]-1\}^{-1}$ is the Bose-Einstein distribution at temperature $T$ and the scattering matrix

### v2

@@ II.2 Two-sublattice model of AFM
$\hbar\omega_{1}({\bf k})$ $=S\sqrt{(A-C_{+}({\bf k}))^{2}-(B-C_{-}({\bf k}))^{2}},$
$\hbar\omega_{2}({\bf k})$ $=S\sqrt{(A+C_{+}({\bf k}))^{2}-(B+C_{-}({\bf k}))^{2}},$ (21)
[…]
${\cal T}({\bf k})=\frac{\sqrt{2}}{2}\begin{pmatrix}-{\cal T}_{1,{\bf k}}&{\cal T}_{2,{\bf k}}&{\cal T}_{1,{\bf k}}^{\prime}&-{\cal T}_{2,{\bf k}}^{\prime}\\ {\cal T}_{1,{\bf k}}&{\cal T}_{2,{\bf k}}&-{\cal T}_{1,{\bf k}}^{\prime}&-{\cal T}_{2,{\bf k}}^{\prime}\\ {\cal T}_{1,{\bf k}}^{\prime}&-{\cal T}_{2,{\bf k}}^{\prime}&-{\cal T}_{1,{\bf k}}&{\cal T}_{2,{\bf k}}\\ -{\cal T}_{1,{\bf k}}^{\prime}&-{\cal T}_{2,{\bf k}}^{\prime}&{\cal T}_{1,{\bf k}}&{\cal T}_{2,{\bf k}}\end{pmatrix}$, (22)
where
${\cal T}_{1,{\bf k}}=\cosh({\varphi}/{2})=(1/\sqrt{2})\sqrt{[A-C_{+}({\bf k})]S/\hbar\omega_{1,{\bf k}}+1},$
${\cal T}_{1,{\bf k}}^{\prime}=\sinh({\varphi}/{2})=(1/\sqrt{2})\sqrt{[A-C_{+}({\bf k})]S/\hbar\omega_{1,{\bf k}}-1},$
${\cal T}_{2,{\bf k}}=\cosh({\phi}/{2})=(1/\sqrt{2})\sqrt{[A+C_{+}({\bf k})]S/\hbar\omega_{2,{\bf k}}+1},$
${\cal T}_{2,{\bf k}}^{\prime}=\sinh({\phi}/{2})=-(1/\sqrt{2})\sqrt{[A+C_{+}({\bf k})]S/\hbar\omega_{2,{\bf k}}-1}.$
Figure 4(b) and (c) plot the spin precession of the two modes: the precession of ${\bf S}_{1}$ and ${\bf S}_{2}$ is out of phase in mode “11”, whereas in mode “22” it is in phase. Both modes have vanished polarization along the Néel vector $\bf n$}. Figure 4(d) addresses the band structures $\omega_{1,2}({\bf k})$ at $\mu_{0}H=2$ T, calculated with $J=4$ meV and $K_{z}=-0.01$ meV; the inset shows enhanced average frequency difference $|\delta\omega|=|\omega_{1}-\omega_{2}|$ by magnetic fields.
[…]
@@
According to Eqs. (9) and (22), the total spin operator $\hat{S}_{z}=({1}/{2})\sum_{{\bf k}}\hat{\Psi}^{\dagger}_{{\bf k}}{\cal S}_{z}({\bf k})\hat{\Psi}_{{\bf k}}$ along the Néel vector with the field operator of magnons $\hat{\Psi}_{{\bf k}}=(\hat{b}_{1,{\bf k}},\hat{b}_{2,{\bf k}},\hat{b}^{\dagger}_{1,-{\bf k}},\hat{b}^{\dagger}_{2,-{\bf k}})^{T}$, in which the spin matrix
${\cal S}_{z}({\bf k})=\begin{pmatrix}0&{\cal A}_{{\bf k}}&0&{\cal B}_{{\bf k}}\\ {\cal A}_{{\bf k}}&0&{\cal B}_{{\bf k}}&0\\ 0&{\cal B}_{{\bf k}}&0&{\cal A}_{{\bf k}}\\ {\cal B}_{{\bf k}}&0&{\cal A}_{{\bf k}}&0\end{pmatrix}$ (23)
with ${\cal A}_{{\bf k}}=({\cal T}_{1,{\bf k}}{\cal T}_{2,{\bf k}}+{\cal T}_{1,{\bf k}}^{\prime}{\cal T}_{2,{\bf k}}^{\prime})\cos\theta$ and ${\cal B}_{{\bf k}}=-({\cal T}_{1,{\bf k}}{\cal T}_{2,{\bf k}}^{\prime}+{\cal T}_{2,{\bf k}}{\cal T}_{1,{\bf k}}^{\prime})\cos\theta$. Accordingly, the total spin operator along the Néel vector
$\hat{S}_{z}=\sum_{{\bf k}}\left({\cal A}_{\bf k}\hat{b}_{1,{\bf k}}^{\dagger}\hat{b}_{2,{\bf k}}+{\cal B}_{\bf k}\hat{b}_{1,{\bf k}}^{\dagger}\hat{b}^{\dagger}_{2,{-\bf k}}\right)+{\rm H.c.}.$ (24)
The population of both magnon modes carries no spin along the Néel vector; the spin is solely carried by the correlations between the two modes.
The velocity of magnon ${\boldsymbol{\cal V}}_{{\bf k}}={\rm diag}({\bf v}_{1,{\bf k}},{\bf v}_{2,{\bf k}},-{\bf v}_{1,-{\bf k}},-{\bf v}_{2,-{\bf k}})$, where ${\bf v}_{n,{\bf k}}=\partial\omega_{n,{\bf k}}/\partial{\bf k}$. According to Eq. (18a), we obtain the uniform spin-current density operator with the spin polarization along the Néel vector $\hat{\bf z}-$direction:
$\hat{J}_{0z}$ $=\frac{1}{2V}\sum_{{\bf k}}({\cal A}_{\bf k}({\bf v}_{1,{\bf k}}+{\bf v}_{2,{\bf k}})\hat{b}_{1,{\bf k}}^{\dagger}\hat{b}_{2,{\bf k}}$
$+{\cal B}_{\bf k}({\bf v}_{1,{\bf k}}+{\bf v}_{2,-{\bf k}})\hat{b}_{1,{\bf k}}^{\dagger}\hat{b}^{\dagger}_{2,{-\bf k}})+{\rm H.c.}.$ (25)
Since ${\cal H}_{\bf k}={\cal H}_{-\bf k}$, ${\bf v}_{2,-\bf k}=-{\bf v}_{2,\bf k}$, such that with ${\bf v}_{1,\bf k}\approx{\bf v}_{2,\bf k}$ only the first term in Eq. (25) contributes, implying that the magnon correlation carries the magnon spin current but not the magnon population. Furthermore, due to the breaking of the $U(1)$ symmetry in canted AFMs, $S_{z}$ is no longer a conserved quantity. The intrinsic torque by Eq. (18b)
$\hat{T}_{0z}$ $=\frac{1}{V}\sum_{{\bf k}}(i{\cal A}_{\bf k}(\omega_{1,{\bf k}}-\omega_{2,{\bf k}})\hat{b}_{1,{\bf k}}^{\dagger}\hat{b}_{2,{\bf k}}$
$+i{\cal B}_{\bf k}(\omega_{1,{\bf k}}+\omega_{2,-{\bf k}})\hat{b}_{1,{\bf k}}^{\dagger}\hat{b}^{\dagger}_{2,{-\bf k}})+{\rm H.c.}$ (26)
is also carried by the magnon correlation.
@@
A solution ${\cal A}_{\bf k}\approx 1$ and ${\cal B}_{\bf k}\approx 0$ when the canted angle is small with $|\gamma|\hbar\mu_{0}H\ll 8JS$ is sufficient to analyze the mechanism. The spin current and torque for spin polarization along $\bf n$} are also solely carried by magnon correlations. Indeed, according to Eqs. (18a) and (18b),
@@
without contribution from the magnon population.
III Electrical injection of magnon correlation
III.1 General magnetic structures
@@
We examine the electrical spin injection from heavy metals to noncollinear magnets in a bilayer structure as illustrated in Fig. 1(b). When a charge current is applied along the $\hat{\bf y}-$direction in the metals, the injected spin current along $\hat{\bf x}$ generates a spin accumulation $\mu_{\uparrow}\neq\mu_{\downarrow}$ polarized along the Néel vector $\hat{\bf z}$ at the interface through the spin Hall effect. This accumulated spin of electrons then excites magnon correlations in the underlying AFM as in Fig. 1(b) via the $s$-$d$ exchange interaction between local spins and conduction electrons at the interface, governed by the Hamiltonian [60, 61, 62, 63]
$\hat{H}_{\rm int}=\sum_{i\in{\rm int}}^{N_{\rm int}}\sum_{l}^{\cal N}{\cal J}_{l}\hat{\bf s}_{i}\cdot\hat{\bf S}_{li},$ (28)
where the summation is performed over the magnetic unit cells “ii”, with $N_{\rm int}$ being their total number in the interfacial plane, and the spin sublattices “ll”. ${\cal J}_{l}$ represents the local exchange coupling, which typically depends on the AFM sublattices. Equal or unequal coupling with different sublattices corresponds to compensated and uncompensated interfaces, as observed experimentally [64, 38, 45].
The electrons in conductors are described in the lattice model with $M$ lattices and the lattice constant $a$. The electron spin at the lattice
$\hat{\bf s}_{i}=\sum_{\alpha,\beta}{\boldsymbol{\sigma}}_{\alpha\beta}\hat{c}^{\dagger}_{\alpha,i}\hat{c}_{\beta,i}=\frac{1}{M}\sum_{\alpha,\beta}\sum_{{\bf k},{\bf k^{\prime}}}{\boldsymbol{\sigma}}_{\alpha\beta}\hat{c}^{\dagger}_{\alpha,{\bf k}-{\bf k^{\prime}}}\hat{c}_{\beta,{\bf k}}e^{i{\bf k^{\prime}}\cdot{\bf r}_{i}},$ (29)
where $\boldsymbol{\sigma}$} is the Pauli matrices and $\hat{c}_{\beta,i}=(1/\sqrt{M})\sum_{{\bf k}}\hat{c}_{{\bf k},\beta}e^{i{{\bf k}}\cdot{\bf r}_{i}}$ is the annihilation operator of electrons. On the other hand, the electron spin density $\hat{\bf s}({\bf r})=\sum_{\alpha,\beta}{\boldsymbol{\sigma}}_{\alpha\beta}\hat{c}^{\dagger}_{\alpha}({\bf r})\hat{c}_{\beta}({\bf r})$. With $\hat{c}_{\beta}({\bf r})=(1/\sqrt{V})\sum_{{\bf k}}\hat{c}_{{{\bf k}},\beta}e^{i{{\bf k}}\cdot{\bf r}}$, where $V=Ma^{3}=N_{\rm int}a^{2}d$ is the volume of the conductors of thickness $d$, we obtain
$\hat{\bf s}({\bf r})=\frac{1}{Ma^{3}}\sum_{\alpha,\beta}\sum_{{\bf k},{\bf k^{\prime}}}{\boldsymbol{\sigma}}_{\alpha\beta}\hat{c}^{\dagger}_{\alpha,{{\bf k}-{\bf k^{\prime}}}}\hat{c}_{\beta,{{\bf k}}}e^{i{{\bf k^{\prime}}}\cdot{\bf r}}.$ (30)
By comparing Eqs. (29) and (30), we find $\hat{\bf s}_{i}=a^{3}\int d{\bf r}\delta({\bf r}-{\bf r}_{i})\hat{\bf s}({\bf r})=a^{3}\hat{\bf s}({\bf r}_{i})$. In terms of the spin density operator $\hat{\bf s}({\bf r})$,
$\hat{H}_{\rm int}=\sum_{i\in{\rm int}}^{N_{\rm int}}\sum_{l}^{\cal N}{\cal J}_{l}a^{3}\int d{\bf r}\delta({\bf r}-{\bf r}_{i})\hat{\bf s}({\bf r})\cdot\hat{\bf S}_{li}.$ (31)
With the local spin operator $\hat{\bf S}_{li}$ in Eq. (4) and $\hat{a}_{li}=(1/\sqrt{N_{\rm int}})\sum_{\bf q}\hat{a}_{l,{\bf q}}e^{i{\bf q}\cdot{\bf r}_{i}}$, the Hamiltonian (31) becomes
$\hat{H}_{\rm int}$ $=\frac{a}{d}\sqrt{\frac{S}{2N_{\rm int}}}\sum_{{\bf k},{\bf k}^{\prime},{\bf q}}\sum_{l=1}^{\cal N}\sum_{\alpha\beta}{\cal J}_{l}{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\xi}}_{l}^{\ast}\hat{c}^{\dagger}_{{{\bf k}^{\prime}},\alpha}\hat{c}_{{{\bf k}},\beta}\hat{a}_{l,{\bf q}}\delta_{{{\bf k}}+{\bf q},{{\bf k}^{\prime}}}$
$+{\rm H.c.}.$ (32)
The bosonic operator $\hat{a}$ is expressed in terms of the magnon operator $\hat{b}$ by the Bogoliubov transformation
$\hat{a}_{l,{\bf q}}=\sum_{\zeta=1}^{\cal N}\left({\cal T}_{l,\zeta}({\bf q})\hat{b}_{\zeta,{\bf q}}+{\cal T}_{l,\zeta+{\cal N}}({\bf q})\hat{b}^{\dagger}_{\zeta,{\bf-q}}\right),$ (33)
where ${\cal T}({\bf q})$ is the Bogoliubov transformation matrix. Substituting Eq. (33) into (32) leads to
$\hat{H}_{\rm int}$ $=\frac{a}{d}\sqrt{\frac{S}{2N_{\rm int}}}\sum_{{\bf k},{\bf k}^{\prime},{\bf q}}\sum_{\zeta=1}^{\cal N}\sum_{\alpha\beta}{\boldsymbol{\sigma}}_{\alpha\beta}\cdot{\boldsymbol{\chi}}^{(\zeta)}_{\bf q}\hat{c}_{{{\bf k}^{\prime}},\alpha}^{\dagger}\hat{c}_{{{\bf k}},\beta}\hat{b}_{\zeta,{\bf q}}\delta_{{{\bf k}}+{\bf q},{\bf k}^{\prime}}$
$+{\rm H.c.},$ (34)
where the vector
${\boldsymbol{\chi}}^{(\zeta)}_{\bf q}=\sum_{l=1}^{\cal N}{\cal J}_{l}\left({\boldsymbol{\xi}_{l}^{\ast}}{\cal T}_{l,\zeta}({\bf q})+{\boldsymbol{\xi}}_{l}{\cal T}^{\ast}_{l,\zeta+\cal N}({\bf q})\right).$ (35)
III.2 Two-sublattice model of AFM
@@
For magnons in the two-sublattice model with a lattice constant $a$ [Fig. 4(a)] interacting with the electrons in a heavy metal of thickness $d$ [60, 61, 62],
@@
$\times\hat{c}_{{{\bf k}^{\prime}},\alpha}^{\dagger}\hat{c}_{{{\bf k}},\beta}\hat{b}_{\zeta,{\bf q}}\delta_{{{\bf k}}+{\bf q},{\bf k}^{\prime}}+{\rm H.c.},$ (36)
@@
$|\boldsymbol{\sigma}_{\alpha\beta}\cdot\boldsymbol{\chi}^{(\zeta)}_{\bf q}|^{2}$ is responsible for the excitation of magnon population in the $\zeta$-mode; $(\boldsymbol{\sigma}_{\alpha\beta}\cdot\boldsymbol{\chi}^{(1)}_{\bf q})(\boldsymbol{\sigma}_{\beta\alpha}\cdot\boldsymbol{\chi}^{(2)\ast}_{\bf q})$ denotes the mixing excitation of both modes in a net spin-flip process, which creates a magnon correlation as in Fig. 1(b). The latter process vanishes in collinear AFMs. Since the magnon population carries no spin along $\hat{\bf z}$, the electron spin is not conserved in the spin-flip scattering (36) since $\boldsymbol{\chi}^{(1,2)}_{\bf q}\nparallel\hat{\bf z}$. When the canted angle $\theta$ is small, $\chi_{z}\rightarrow 0$ such that the interfacial spin-conserving scattering $|\boldsymbol{\sigma}_{\alpha\alpha}\cdot\boldsymbol{\chi}^{(\zeta)}_{\bf q}|^{2}\rightarrow 0$. Consequently, the excitation of magnon correlation is dominated by electron spin-flip scattering.
III.3 Electrical spin injection and dephasing
To address how magnon correlation is injected, how it relaxes, and how it transports, we establish the quantum kinetic equation for magnons. In the density matrix ${\rho}({\bf q})=\langle\hat{\rho}_{\bf q}\rangle=\left\langle(\hat{\Psi}^{\dagger}_{\bf q})^{T}\hat{\Psi}^{T}_{\bf q}\right\rangle$ of magnons with $\langle\cdots\rangle$ denoting the ensemble average, the diagonal elements represent the magnon population, while the off-diagonal elements denote the correlation between different magnon modes. The quantum kinetic equation $\partial_{t}{\rho}_{\bf q}=\partial_{t}{\rho}_{\bf q}|_{\rm coh}+\partial_{t}{\rho}_{\bf q}|_{\rm inj}+\partial_{t}{\rho}_{\bf q}|_{\rm rex}$ contains three contributions (refer to Appendix A for the derivation). The coherent dynamics
$\partial_{t}{\rho}_{\bf q}|_{\rm coh}=(1/i\hbar)\left({\rho}_{\bf q}\sigma_{3}{\cal E}_{\bf q}-{\cal E}_{\bf q}\sigma_{3}{\rho}_{\bf q}\right),$ (38)
induced by the bare Hamiltonian $\hat{H}_{m}=\sum_{{\bf q},\zeta}\hbar\omega_{\zeta,{\bf q}}\hat{b}_{\zeta,{\bf q}}^{\dagger}\hat{b}_{\zeta,{\bf q}}$, originates from the intrinsic spin torque (18b). The exchange of spin with metals due to $\hat{H}_{\rm int}$ is governed by
$\partial_{t}{\rho}_{\bf q}|_{\rm inj}$ $=\frac{\pi}{\hbar}\frac{S}{2N_{\rm int}}\frac{a^{2}}{d^{2}}\sum_{{\bf k},{\bf k^{\prime}}}\sum_{\alpha,\beta}[f_{{\bf k^{\prime}},\alpha}(1-f_{{\bf k},\beta})(\sigma_{3}+{\rho}_{\bf q})$
$-f_{{\bf k},\beta}(1-f_{{\bf k^{\prime}},\alpha})\rho_{\bf q}]{\cal M}_{\alpha\beta}({\bf k},{\bf k}^{\prime})\delta_{{\bf k}+{\bf q},{\bf k}^{\prime}}+{\rm H.c.},$ (39)
which may be intuitively interpreted by Fermi’s Golden Rule: $f_{{\bf k^{\prime}},\alpha}(1-f_{{\bf k},\beta})(\sigma_{3}+\rho_{{\bf q}})$ represents the magnon excitation process and $(1-f_{{\bf k^{\prime}},\alpha})f_{{\bf k},\beta}\rho_{{\bf q}}$ represents the magnon absorption process. The scattering potentials are included in the $4\times 4$ matrix
${\cal M}_{\alpha\beta}({\bf k},{\bf k^{\prime}})=\begin{pmatrix}{A}_{2\times 2}({\bf k},{\bf k^{\prime}})&{C}_{2\times 2}({\bf k},{\bf k^{\prime}})\\ {D}_{2\times 2}({\bf k},{\bf k^{\prime}})&{B}_{2\times 2}({\bf k},{\bf k^{\prime}})\end{pmatrix}$. (40)
In the blocks